OpenStudy (anonymous):

Find the value of k such that when 2x^3+9x^2+kx-15 is divided by x+5, the remainder is 0.

OpenStudy (anonymous):

If I can remember correctly.. My solution is: x + 5 = 0 x = -5 substitute -5 to x then solve for k. Use the remainder 0 in the left hand/right hand side of the equation. 2x^3+9x^2+kx-15 = 0 (The remainder) 2(-5)^3 + 9(-5)^2 + k (-5) - 15 = 0 -250 + 225 - 15 - 5k = 0 85 - 5k = 0 -5k = -85 5k = 85 k = 17